OpenStudy (anonymous):

Ive gone through with someone on parts one and two on the question below, but I am stuck on the last part : "are the graphs of: inverse tan(x) + inverse tan (y) = -pi/4 AND y=((x+1)/(x-1)) from above the same? EXPLAIN" Can anyone help on that one? I was thinking that the graphs would be the same since they are the same equation rewritten? But I was second guessing that because maybe they would look different rewritten in different forms? ORIGINAL QUESTION BELOW: show that the solution of: dy/dx = -((1+y^2)/(1+x^2)) y(0) = -1 is the inverse tan (x) + inverse tangent (y) = -pi/4 AND show that the above can be written : y= ((x+1)/(x-1)) And are the graphs of: inverse tan(x) + inverse tan (y) = -pi/4 AND y=((x+1)/(x-1)) from above the same? Explain.

OpenStudy (freckles):

have you implicitly found the derivative of arctan(x)+arctan(y)=-pi/4

OpenStudy (freckles):

Also is y(0)=-1 when arctan(x)+arctan(y)=-pi/4

OpenStudy (freckles):

in other words does this hold: arctan(0)+arctan(-1)=-pi/4 ?

OpenStudy (anonymous):

Yes I think i answered part one correctly, that inverse tanx + inverse tany = -pi/4, I am having a lot of trouble on the next two parts more so I guess

OpenStudy (anonymous):

I have been trying to work backwards from the second derivative to the original function but it is not working out, am I on the right track?

OpenStudy (freckles):

why are you finding second derivative?

OpenStudy (freckles):

I would think all you need to do for the first part is confirm arctan(0)+arctan(-1)=-pi/4 and the implicity find the deirvative of arctan(x)+arctan(y)=-pi/4 and solve for dy/dx . For the second part it sounds like we need to solve arctan(x)+arctan(y)=-pi/4 for y

OpenStudy (freckles):

\[\arctan(y)=\frac{-\pi}{4}-\arctan(x) \\ y=\tan(\frac{-\pi}{4}-\arctan(x)) \\ \text{ use difference identity for tan }\]

OpenStudy (anonymous):

I meant first derivative, looking over it again, I tried to take a first derivative of y=(x+1)/(x-1) but I am not coming up with the original dy/dx function

OpenStudy (anonymous):

I don't know if that even makes sense, I am so confused.

OpenStudy (freckles):

are you trying to do the second write now ?

OpenStudy (freckles):

second part?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

then this is the way I would proceed \[\arctan(y)=\frac{-\pi}{4}-\arctan(x) \\ y=\tan(\frac{-\pi}{4}-\arctan(x)) \\ \text{ use difference identity for \tan } \]

OpenStudy (freckles):

@Emilyjones85 is that going okay for you?

OpenStudy (anonymous):

so that would be (tan(-pi/4) - tan (tan inverse (x)))/(1+tan (-pi/4) tan(inverse tan(x)) ?

OpenStudy (freckles):

yes y= but yes

OpenStudy (anonymous):

yes sorry I forgot that

OpenStudy (anonymous):

so tan(-pi/4) = -1 right?

OpenStudy (freckles):

do you know what tan(-pi/4)=? and tan(arctan(x))=?

OpenStudy (freckles):

and yes on the first one

OpenStudy (freckles):

I will give you a hint tan and arctan are inverse functions

OpenStudy (freckles):

the domain of arctan(x) is actually all real numbers

OpenStudy (freckles):

so tan(arctan(x)) will always equal just x

OpenStudy (freckles):

i might be right back i have to go do something real quick

OpenStudy (anonymous):

right, so its (-1-x)/(1-x) - how does the numerator become x+1 though, am I missing something.

OpenStudy (anonymous):

And on the last part would it just be that they are equal because they are the same thing?

OpenStudy (freckles):

\[y=\frac{-1-x}{1-x}=\frac{-1(1+x)}{1-x}=\frac{1+x}{(-1)(1-x)}\]

OpenStudy (freckles):

since -1/1=1/-1

OpenStudy (anonymous):

Ok got it, sorry I had to get off last night. So that all makes sense now. And Like i was saying, the graphs would be the same, right? Because they turn out to be the same functions, right?